OpenStudy (anonymous):

Differentiate. Answer with all positive exponents. 1) h(r)= (r^2+4r-6)/(9r+8)

OpenStudy (anonymous):

Do you know quotient rule?

OpenStudy (anonymous):

Isnt that rewriting the equation like this (r^2+4r-6)x(9r+8)+(r^2+4r-6)x(9r+8)?

OpenStudy (anonymous):

But taking the derivative once on each side right?

OpenStudy (anonymous):

\[\frac{d}{dx}\frac{f(x)}{g(x)}=\frac{f'(x)g(x)-f(x)g'(x)}{g(x)^2}\]

OpenStudy (anonymous):

Ohh wow thanks. That's helpful to finally see something like that.

OpenStudy (anonymous):

(r^2+4r-6)'x(9r+8)-(r^2+4r-6)x(9r+8)' /(9r+8)^2 is that right?

OpenStudy (anonymous):

Then you just need to do the differentiation and simplify (optional).

OpenStudy (anonymous):

Could I work it out and have you try it?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

9r^2+16r+86/72r^2+126r+64?

OpenStudy (anonymous):

\[\huge\frac{ (2r+4)(9r+8)-9(r^2+4r-6) }{ (9r+8)^2 }\]

OpenStudy (anonymous):

The numerator is correct, but there is some calculation mistakes in the denominator.

OpenStudy (anonymous):

Thanks deena123

OpenStudy (anonymous):

\[(9r^2+8)=81r^2+144r+64\]

OpenStudy (anonymous):

but the denominator doesn't need to be multiplied right agostino?

OpenStudy (anonymous):

Typo lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[(9r+8)\]

OpenStudy (anonymous):

could i leave it as (9r+8)^2?

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yeah, you can.

OpenStudy (anonymous):

Sometimes it is better to leave it there as you may cross it out with the numerator.

OpenStudy (anonymous):

Alright, thanks for the help you guys, I'm new to this site so is there a way I can honor you guys?

OpenStudy (anonymous):

Participate more and be active. :P